Express to a positive exponent then find its value
(2/3)^-2

Answer by stanbon(75887)

Express to a positive exponent then find its value
(2/3)^-2
The negative sign in the exponent means "invert"
--------------
So, (2/3)^-2 = (3/2)^2 = 9/4
------------
Cheers,
Stan H.